# Chart-R1: Chain-of-Thought Supervision and Reinforcement for Advanced Chart Reasoner

### v2

@@ Abstract
Recently, inspired by OpenAI-o1/o3 and Deepseek-R1, the R1-Style method based on reinforcement learning fine-tuning has received widespread attention from the community. Previous R1-Style methods mainly focus on mathematical reasoning and code intelligence. It is of great research significance to verify their advantages on more general multimodal data. Chart is an important multimodal data type with rich information, which brings important research challenges in complex reasoning. In this work, we introduce Chart-R1, a chart-domain vision-language model with reinforcement learning fine-tuning to enable complex chart reasoning. To support Chart-R1, we first propose a novel programmatic data synthesis technology to generate high-quality step-by-step chart reasoning data covering single- and multi-subcharts, which makes up for the lack of reasoning data in the chart domain. Then we develop a two-stage training strategy: Chart-COT with step-by-step chain-of-thought supervision, and Chart-RFT with numerically sensitive reinforcement fine-tuning. Chart-COT aims to decompose complex chart reasoning tasks into fine-grained, understandable subtasks through step-by-step supervision, which lays a good foundation for improving the reasoning level of reinforcement learning. Chart-RFT utilize the typical group relative policy optimization strategy, in which a relatively soft reward is adopted for numerical response to emphasize the numerical sensitivity in the chart domain. We conduct extensive experiments on open-source benchmarks and self-built chart reasoning dataset (i.e., ChartRQA). Experimental results show that Chart-R1 has significant advantages compared to chart-domain methods, even comparable to open/closed source large-scale models (e.g., GPT-4o, Claude-3.5). Our code and dataset will be available at https://github.com/DocTron-hub/Chart-R1.
@@ 1 Introduction
Recently, inspired by the success of models such as OpenAI’s o1/o3 OpenAI (2025) and DeepSeek-R1 Guo et al. (2025), leveraging Reinforcement Learning (RL) for fine-tuning has garnered significant attention within the research community. Although these methods have shown promise in textual domains like mathematical reasoning, code generation, and multidisciplinary knowledge, transferring these advanced reasoning capabilities to the vision domain presents an open challenge. While recent approaches like Vision-R1 Huang et al. (2025) and VLM-R1 Shen et al. (2025) have successfully leveraged RL to enhance visual perception and grounding, they have primarily focused on simple questions, neglecting tasks that demand deep reasoning capabilities.
@@
Charts, as information-intensive images, are a crucial research area in image understanding and reasoning Wang et al. (2024). Prior works improve chart perception and understanding capacities by supervised fine-tuning (SFT) on datasets augmented with Chain-of-Thought (CoT) or Program-of-Thought (PoT) methods Wei et al. (2022); Chen et al. (2022). A key limitation of SFT is that it causes models to overfit specific reasoning patterns, hindering their generalization abilities. Following the DeepSeek R1, recent methods Jia et al. (2025); Ni et al. (2025) leverage RL to enhance VLM reasoning capabilities. However, the scope of these efforts has been largely limited to visual perception and understanding, rather than the complex reasoning required for deep chart analysis.
@@
Figure 1: Comparison of existing VLM performance on chart understanding and reasoning benchmarks. In the complex chart reasoning task, Qwen2.5-VL-7B generates a wrong thinking process, whereas Chart-R1 thinks and answers correctly.
@@
In this work, we propose Chart-R1, a chart domain VLM that leverages RL to enhance complex reasoning capability. To this end, we introduce two key contributions. First, we propose a novel programmatic synthesis strategy to generate high-quality reasoning data. Second, we introduce an effective two-stage training strategy that significantly enhances reasoning capacity. Specifically, to support Chart-R1 training, we first generate complex chart reasoning data in the programmatic synthesis method. We utilize LLMs to generate the chart plotting code and then use the generated code to formulate complex questions, multi-step CoT reasoning processes, and the final answer. To this end, we construct ChartRQA, a complex reasoning dataset featuring 258k multi-step reasoning samples that cover both single- and multi-chart tasks. To ensure the fidelity of the data in charts, we curate real-world tables from arXiv papers as the data source. ChartReasoner Jia et al. (2025) proposes converting charts into code to generate reasoning data. However, its reliance on a lossy parsing process directly restricts the scope and diversity of the final reasoning data. The training of Chart-R1 is conducted in two stages: Chart-COT with step-by-step chain-of-thought supervision, and Chart-RFT with numerically sensitive reinforcement fine-tuning. During the initial Chart-COT stage, the model is fine-tuned via SFT on step-by-step reasoning data to build its core capability of decomposing complex tasks into fine-grained subtasks. In the Chart-RFT stage, we utilize the group relative policy optimization (GRPO) strategy, where the reward signal is a composite of soft matching and edit distance. This design specifically targets and enhances accuracy for both numerical and string-based answers. Notably, distinct datasets are employed for these two stages, based on our finding that training on the same data impairs the model’s exploration ability during the RL process. Furthermore, we introduce a human-verified benchmark, ChartRQA, to assess the boundaries of complex chart reasoning. Unlike prior works Xia et al. (2024); Wang et al. (2024), its questions feature a higher degree of complexity, requiring multi-step thought processes. The substantial performance drop of existing VLMs on ChartRQA exposes a critical limitation in their chart reasoning capabilities.
In summary, our contributions are as follows:
@@
  • We propose a novel two-stage training strategy, consisting of Chart-COT and Chart-RFT, to enhance chart reasoning in VLMs. Our model, Chart-R1, trained with this strategy, establishes a new SOTA on various chart understanding and reasoning benchmarks.
@@
  • We introduce a programmatic data synthesis strategy that leverages code as a pivotal starting source to generate step-by-step reasoning data. The data source is grounded in real-world tables from arXiv papers, ensuring high fidelity in the resulting charts.
  • We introduce ChartRQA, a comprehensive dataset for complex chart reasoning that includes a human-verified benchmark and a large-scale training dataset. The substantial performance of existing VLMs on the ChartRQA benchmark underscores a critical limitation in their chart reasoning capabilities.
  • We conduct a series of comprehensive experiments to systematically assess the impact of various settings. Our findings provide valuable insights and offer clear guidance for future research in this domain.
@@ 2.1 Chart VLMs
Chart understanding and reasoning are crucial areas of research community that encompass both low-level and high-level tasks Singh et al. (2019); Methani et al. (2020). Recently, many chart-domain models have been proposed to enhance the chart understanding capacity of VLMs Han et al. (2023); Liu et al. (2023). However, prior works have concentrated on descriptive tasks Masry et al. (2024a, b), such as extracting explicit content from charts Masry et al. (2022). In contrast, more recent works focus on leveraging the reasoning capabilities of VLMs to interpret complex and implicit information within the charts. For example, TinyChart Zhang et al. (2024) utilizes a template-based method to generate the Program-of-Thought (PoT) Chen et al. (2022) reasoning data. ChartCoder Zhao et al. (2025b) proposes Snippet-of-Thought to enhance chart-to-code generation. ChartReasoner Jia et al. (2025) utilizes a chart-to-code model to convert chart images into code and generate the reasoning process based on code. However, the generated reasoning data has limitations due to the chart-to-code accuracy Shi et al. (2024); Xu et al. (2024).
@@ 2.2 Long Reasoning VLMs
Recently, with the success of DeepSeek-R1 Guo et al. (2025), many works have attempted to enhance the LLM reasoning ability via rule-based reward and RL Shao et al. (2024). In the vision-language domain, recent works follow the DeepSeek-R1 method to enhance the long-chain reasoning capacity of VLMs Shen et al. (2025); Wang et al. (2025); Qiu et al. (2025). For example, Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025) apply Group Relative Policy Optimization (GRPO) with multimodal reasoning data to enable VLMs for long reasoning. MMEureka Meng et al. (2025b) and R1-Zero Liu et al. (2025) further advance the visual long-term reasoning with improved RL training strategies. Point-RFT Ni et al. (2025) utilizes grounded CoT reasoning for visual understanding, but it just utilize ChartQA for RL which limits the final model reasoning capacity.
@@ 2.3 Chart Understanding and Reasoning
A variety of training datasets and evaluation benchmarks have been developed to improve VLM performance on chart-related tasks Xia et al. (2024); Shi et al. (2024); He et al. (2024); Zhao et al. (2025a); Wu et al. (2025). Previous works generally focus on description tasks, for example, ChartQA Masry et al. (2022), PlotQA Methani et al. (2020) and Chart-to-text Kantharaj et al. (2022) mainly train and evaluate the capacities of the models on extracting information from the chart. While numerous relevant works exist, the challenge in the description tasks is predominantly driven by chart complexity. Recent works such as Charxiv Wang et al. (2024) and CharMuseum Tang et al. (2025) introduce more challenging reasoning tasks, demanding that models think before answering. Unlike descriptive tasks, reasoning tasks present a dual challenge, originating from both the perceptual complexity of the chart and the reasoning depth required by questions.
@@ 3.1 Programatic Data Synthesis
While several CoT datasets for chart reasoning have been proposed, they are largely derivatives of the ChartQA dataset, constructed by augmenting its existing question-answer pairs with generated reasoning processes Zhang et al. (2024); Jia et al. (2025). However, this method is akin to distilling reasoning from SOTA VLMs, which is problematic as the failures of these models on complex tasks inherently limit the quality of the generated data. Generating high-quality CoT reasoning data is a well-recognized challenge, largely because current methods use the final answer’s correctness as the sole supervisory signal. This problem is particularly acute in the domain of complex chart reasoning, as existing models already exhibit significant limitations. Consequently, data generated via this approach inherently suffers from both low quality and limited diversity. Although the recent ChartReasoner method Jia et al. (2025) generates reasoning data by first parsing charts into code, the diversity and quality of generated data is fundamentally limited by the performance of the chart-to-code parser. In contrast, our programmatic data generation strategy reverses this paradigm by utilizing code as a pivotal starting source. First, we prompt a powerful LLMs to generate plotting code. This code then serves as a perfect, high-fidelity foundation from which the LLM subsequently synthesizes question-answer pairs and their complex step-by-step reasoning path.
@@
Reasoning Data Generation With the executable plotting code as a foundation, we prompt LLMs to synthesize a complete reasoning instance, comprising a question, its answer, and a step-by-step reasoning path. To increase the diversity, we separate the plotting code into single- and multi-chart ones and utilize distinct instructions for instance generation. For multi-chart problems in particular, we prompt the LLM to generate questions requiring information to be cross-referenced between sub-charts. The results show that this strategy significantly enhances multi-chart task complexity. Our results show that leveraging code allows LLMs to produce more complex questions and detailed reasoning compared to methods that use chart images alone. We posit that a code-based approach is superior for generating complex chart reasoning as the underlying code provides a lossless textual representation of details while enabling the scalable synthesis of new data independent of existing corpora. We filter out data samples that do not conform to the thinking and answering formats and faulty chart images.
Dataset Construction Using the aforementioned methods, we construct ChartRQA, a comprehensive chart reasoning corpus that includes a large-scale training dataset of 258k instances with reasoning paths as well as a human-verified benchmark. The training dataset is separated into two subsets for our two-stage training strategy, ChartRQA-SFT and ChartRQA-RL, consisting of 228k and 30k samples, respectively. Detailed comparisons about ChartRQA with other chart-domain training set are denoted in Table 1. The benchmark is constructed via a human validation where experts review each sample for question difficulty and answer correctness, subsequently constructing 1,702 high-quality samples (933 single-chart and 769 multi-chart tasks) for evaluation. As detailed in Table 2, we also calculated the average token counts for the questions, reasoning paths, and final answers, broken down by single- and multi-chart problems. The analysis reveals that the components associated with multi-chart problems are significantly longer than those for single-chart problems. Also, the distribution between the train and test sets is balanced. Figure 2 is the showcase of our generated ChartRQA.
Token Avg. Train Test
Single Multi Total Single Multi Total
Question 30.03 39.84 34.03 29.83 39.49 34.19
Thinking Process 196.50 237.38 213.17 196.32 240.94 216.48
Answer 5.98 8.87 7.16 5.96 8.97 7.32
@@ 3.2 Chart-COT
To enhance the chart reasoning capacity, we propose a two-stage training strategy. Utilizing Qwen2.5VL-7B-Instruct as the baseline model, we first SFT it on the step-by-step reasoning data of our proposed ChartRQA-SFT. Specifically, the baseline model first undergoes SFT on our generated step-by-step reasoning data, which serves as the code-starting phase to equip the model with the fundamental capability to decompose complex tasks into fine-grained subtasks. Our ablation studies demonstrate that a preliminary SFT stage on CoT data is critical, as it yields significantly better performance than applying RL from scratch.
We train the model using a standard autoregressive language modelling objective. The loss function is the negative log-likelihood of the target sequence:
$\mathcal{L}(\theta):=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{CoT}}}\sum_{t=1}^{T}\log P\left(y_{t}\mid x,y_{<t};\theta\right),$ (1)
where $(x,y)$ is the query and target response, with the reasoning process.
@@ 3.3 Chart-RFT
After the Chart-COT stage, while the fine-tuned model demonstrates an enhanced ability to decompose complex questions, its performance on out-of-domain (OOD) tasks notably degrades. We hypothesize this is due to a distributional mismatch between ChartRQA-SFT with some simple chart understanding tasks, which harms its generalization ability. To address the degradation in generalization, we subsequently apply reinforcement fine-tuning (RFT) to generalize its reasoning capacity.
Group Relative Policy Optimization Following recent reasoning works Guo et al. (2025), we adapt the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm for RFT. GRPO foregoes the critic model, instead estimating the baseline from group scores, significantly reducing training resources. For each input $(x,y)$, the policy $\pi_{\theta}$ samples a group of $G$ candidate responses $\{o_{i}\}_{i=1}^{G}$.
$\mathcal{J}_{GRPO}(\theta)=$ $\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D}_{\text{CoT}},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid x)\end{subarray}}\biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\biggl{(}\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)}A_{i},$ (2)
$\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)},1-\varepsilon,1+\varepsilon\right)A_{i}\biggr{)}-\beta\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{\mathrm{SFT}})\biggr{]}$
where $\varepsilon$ and $\beta$ are hyperparameters, and $\pi_{\mathrm{SFT}}$, $\pi_{\theta}$, and $\pi_{\theta_{\text{old}}}$ are the model after SFT, the optimized model and the old policy model. The group-normalized advantage for the $i$-th response is:
@@
Reward Design For effective RFT, we follow the DeepSeek-R1 Shao et al. (2024) and adopt a rule-based reward that consists of accuracy and format rewards. We introduce a soft accuracy reward for chart problems, which utilizes distinct functions to evaluate numerical and string-based tasks separately. This allows for a more appropriate assessment based on the expected answer type.
  • Accuracy Reward. We employ distinct reward functions to measure the correctness of model outputs, with each function tailored to the specific answer type. In the case of numerical answers, we adopt the soft matching technique from Point-RFT Ni et al. (2025), allowing for a relative error tolerance of $\pm$5%. For string-based answers, we utilize the edit distance as the reward signal.
  • Format Reward. The format reward is determined using a grammar-level regex parser. This parser validates the structural integrity of the output by checking two conditions: (1) that the reasoning process is correctly enclosed within <think> and </think> tags, and (2) that the final answer can be extracted from the designated answer tag <answer> and </answer>.
@@
Figure 2: Showcases of our proposed ChartRQA dataset. The ChartRQA include single- and multi-chart images with complex questions that need step-by-step thinking processes to answer.

### v3

@@ Abstract
Chart reasoning presents unique challenges due to its inherent complexity—requiring precise numerical comprehension, multi-level visual understanding, and logical inference across interconnected data elements. Existing vision-language models often struggle with such reasoning tasks, particularly when handling multi-subchart scenarios and numerical sensitivity. To address these challenges, we introduce Chart-R1, a chart-domain vision-language model that leverages reinforcement fine-tuning for advanced chart reasoning. We first propose a programmatic data synthesis approach to generate high-quality step-by-step reasoning data with verifiable answer formats, covering diverse chart types and complexity levels. Our two-stage training strategy includes: (1) Chart-COT, which decomposes complex reasoning into interpretable subtasks through chain-of-thought supervision, and (2) Chart-RFT, which employs group relative policy optimization with numerically sensitive rewards tailored for chart-specific reasoning. Experiments on open-source benchmarks and our proposed ChartRQA dataset demonstrate that Chart-R1 significantly outperforms existing chart-domain methods and rivals large-scale open/closed-source models.
Chart-R1: Chain-of-Thought Supervision and Reinforcement
for Advanced Chart Reasoner
Lei Chen   Xuanle Zhao   Zhixiong Zeng   Jing Huang   Yufeng Zhong   Lin Ma Meituan leichen1997@outlook.com  zengzhixiong@meituan.com  forest.linma@gmail.com
footnotetext: Project leader. Corresponding author.
@@ 1 Introduction
Recently, inspired by the success of models such as OpenAI’s o1/o3 OpenAI (2025b) and DeepSeek-R1 Guo et al. (2025), leveraging Reinforcement Learning (RL) for fine-tuning has garnered significant attention within the research community. Although these methods have shown promise in textual domains like mathematical reasoning, code generation, and multidisciplinary knowledge, transferring these advanced reasoning capabilities to the vision domain presents an open challenge. While recent approaches like Vision-R1 Huang et al. (2025) and VLM-R1 Shen et al. (2025) have successfully leveraged RL to enhance visual perception and grounding, they have primarily focused on simple questions, neglecting tasks that demand deep reasoning capabilities.
@@
Charts, as information-intensive images, are a crucial research area in image understanding and reasoning Wang et al. (2024). Chart reasoning requires decomposing complex questions into multiple interpretable steps involving precise numerical comprehension, visual parsing, and logical inference. As illustrated in Figure 5, questions with fewer reasoning steps are relatively straightforward, but the difficulty increases substantially when problems require multi-step decomposition. Prior works attempt to improve chart reasoning through supervised fine-tuning (SFT) on datasets augmented with Chain-of-Thought (CoT) or Program-of-Thought (PoT) methods Wei et al. (2022); Chen et al. (2022). However, SFT causes models to overfit specific reasoning patterns, hindering their generalization abilities. Recent methods Jia et al. (2025); Ni et al. (2025) leverage RL to enhance VLM reasoning capabilities, yet their focus remains primarily on perception and understanding rather than the multi-step reasoning required for complex chart analysis.
@@
Figure 1: Comparison of Qwen2.5-VL-7B and Chart-R1 on chart understanding and reasoning benchmarks. In the complex chart reasoning task, Qwen2.5-VL-7B generates a wrong thinking process, whereas Chart-R1 thinks and answers correctly.
@@
In this work, we propose Chart-R1, a chart-domain VLM that leverages RL to enhance advanced multi-step reasoning capability, which achieves superior performance as shown in Figure 1. To support Chart-R1, we introduce two key contributions. First, we propose a programmatic synthesis strategy to generate high-quality step-by-step reasoning data with verifiable answer formats, covering diverse chart types and complexity levels. Specifically, we utilize LLMs to generate the chart plotting code from real-world tables curated from arXiv papers, ensuring data fidelity in the resulting charts. The generated code then serves as a foundation for formulating complex questions, multi-step CoT reasoning processes, and verifiable answers, resulting in ChartRQA—a comprehensive dataset featuring 258k multi-step reasoning samples covering both single- and multi-chart scenarios. Second, we develop an effective two-stage training strategy: (1) Chart-COT with step-by-step chain-of-thought supervision, and (2) Chart-RFT with numerically sensitive reinforcement fine-tuning. During the Chart-COT stage, the model learns to decompose complex tasks into fine-grained, interpretable subtasks through supervised learning on step-by-step reasoning data. As shown in Figure 1, this explicit decomposition capability is crucial for handling complex scenarios that require multi-step reasoning. In the Chart-RFT stage, we employ group relative policy optimization (GRPO) with a composite reward signal of soft matching and edit distance to enhance accuracy for both numerical and string-based answers. We utilize distinct datasets for these two stages, as our findings show that training on the same data impairs the model’s exploration ability during the RL process. Furthermore, we introduce ChartRQA, a human-verified benchmark designed to probe the limits of complex chart reasoning. In contrast to prior works Wang et al. (2024), its questions demand higher complexity and multi-step thought processes. The substantial performance drop of existing VLMs on ChartRQA reveals a critical gap in their reasoning capabilities. In summary, our contributions are as follows:
@@
  • To enhance chart reasoning in VLMs, we propose a novel two-stage training strategy consisting of Chart-COT and Chart-RFT. The resulting model, Chart-R1, sets a new state-of-the-art across various chart understanding and reasoning benchmarks.
@@
  • We introduce a programmatic data synthesis strategy that leverages code generation from real-world tables to produce step-by-step reasoning data with verifiable answers, ensuring both data fidelity and reasoning quality.
  • We present ChartRQA, a comprehensive dataset with a human-verified benchmark and 258k training samples. Existing VLMs show significant performance gaps on this benchmark, revealing critical reasoning limitations.
@@ 2.1 Chart VLMs
Chart understanding and reasoning are crucial areas of research community that encompass both low-level and high-level tasks Singh et al. (2019); Methani et al. (2020). Recently, many chart-domain models have been proposed to enhance the chart understanding capacity of VLMs Han et al. (2023); Liu et al. (2023). However, prior works concentrate on descriptive tasks Masry et al. (2024a, b), such as extracting explicit content from charts Masry et al. (2022). More recent works focus on leveraging reasoning capabilities to interpret complex and implicit information. For example, TinyChart Zhang et al. (2024) utilizes a template-based method to generate Program-of-Thought (PoT) Chen et al. (2022) reasoning data. ChartCoder Zhao et al. (2025b) proposes Snippet-of-Thought for chart-to-code generation. ChartReasoner Jia et al. (2025) utilizes a chart-to-code model to convert images into code and generate reasoning processes. However, the generated reasoning data has limitations due to chart-to-code accuracy Shi et al. (2024); Xu et al. (2024).
@@ 2.2 Long Reasoning VLMs
Recently, with the success of DeepSeek-R1 Guo et al. (2025), many works attempt to enhance LLM reasoning ability via rule-based reward and RL Shao et al. (2024). In the vision-language domain, recent works follow the DeepSeek-R1 method to enhance the long-chain reasoning capacity of VLMs Shen et al. (2025); Wang et al. (2025b); Qiu et al. (2025). For example, Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025a) apply Group Relative Policy Optimization (GRPO) with multimodal reasoning data to enable long reasoning. MMEureka Meng et al. (2025a) and R1-Zero Liu et al. (2025) further advance visual long-term reasoning with improved RL training strategies. Point-RFT Ni et al. (2025) uses grounded CoT for visual understanding, but it only utilize ChartQA for RL which limits the model’s reasoning capacity.
Figure 2: ChartRQA dataset pipeline via programmatic data synthesis. The pipeline consists of plotting code generation and reasoning data generation, producing ChartRQA-SFT, ChartRQA-RL, and the ChartRQA benchmark. Examples show single-chart and multi-chart scenarios requiring multi-step reasoning.
@@ 2.3 Chart Understanding and Reasoning
A variety of training datasets and evaluation benchmarks have been developed to improve VLM performance on chart-related tasks Xia et al. (2024); Shi et al. (2024); He et al. (2024); Zhao et al. (2025a); Wu et al. (2025). Previous works generally focus on description tasks. For example, ChartQA Masry et al. (2022), PlotQA Methani et al. (2020) and Chart-to-text Kantharaj et al. (2022) mainly train and evaluate models’ capacity on extracting information from charts, where the challenge is predominantly driven by chart complexity. Recent works such as Charxiv Wang et al. (2024) and CharMuseum Tang et al. (2025a) introduce more challenging reasoning tasks that require models to think before answering. Unlike descriptive tasks, reasoning tasks present a dual challenge from both the perceptual complexity of charts and the reasoning depth required by questions.
@@ 3 Method
Token Avg. Train Test
Single Multi Single Multi
Question 30.03 39.84 29.83 39.49
Thinking Process 196.50 237.38 196.32 240.94
Answer 5.98 8.87 5.96 8.97
@@ 3.1 Programatic Data Synthesis
While several CoT datasets for chart reasoning have been proposed, they are largely derivatives of the ChartQA dataset, constructed by augmenting its existing question-answer pairs with generated reasoning processes Zhang et al. (2024); Jia et al. (2025). However, this method is like distilling reasoning from top VLMs, so it naturally inherits their limitations and errors on complex tasks. The reliance on final answer correctness as the only supervisory signal makes generating high-quality CoT reasoning data a significant challenge. This issue is amplified in complex chart reasoning, where the struggles of even top models inherently lead to low-quality, undiverse data. Although the recent ChartReasoner method Jia et al. (2025) generates reasoning data by first parsing charts into code, the diversity and quality of generated data are fundamentally limited by the performance of the chart-to-code parser. In contrast, our programmatic data generation strategy reverses this paradigm by utilizing code as a pivotal starting source. First, we prompt a powerful LLM to generate plotting code. This code then serves as a perfect, high-fidelity foundation from which a VLM subsequently synthesizes question-answer pairs and their complex step-by-step reasoning path. An overview of our data synthesis pipeline is shown in Figure 2.
@@
Reasoning Data Generation With the executable plotting code as a foundation, we prompt LLMs to synthesize a complete reasoning instance, comprising a question, its answer, and a step-by-step reasoning path. To enhance diversity, we categorize the plotting code into single- and multi-chart types and apply distinct generation instructions for each. For multi-chart problems, we instruct the LLM to generate questions that necessitate cross-referencing information between sub-charts. The generated data show that this strategy significantly enhances multi-chart task complexity. Our results show that leveraging code allows LLMs to produce more complex questions and detailed reasoning compared to methods that use chart images alone. We posit that a code-based approach is superior for generating complex chart reasoning as the underlying code provides a lossless textual representation while enabling the scalable synthesis of data independent of existing corpora.
Dataset Construction Using the aforementioned methods, we construct the ChartRQA dataset, which includes a large-scale training dataset of 258k instances with reasoning paths as well as a human-verified benchmark. The training dataset is separated into two subsets for our two-stage training strategy, ChartRQA-SFT and ChartRQA-RL, consisting of 228k and 30k samples, respectively. Detailed comparisons about ChartRQA with other chart-domain training sets are denoted in Table 1. The benchmark is constructed via a human validation where experts review each sample for question difficulty and answer correctness, subsequently constructing 1,702 high-quality samples (933 single-chart and 769 multi-chart tasks) for evaluation. As detailed in Table 2, we also calculated the average token counts for the questions, reasoning paths, and final answers, broken down by single- and multi-chart problems. The analysis reveals that the components associated with multi-chart problems are significantly longer than those for single-chart problems.
@@ 3.2 Chart-COT
To enhance the chart reasoning capacity, we propose a two-stage training strategy. Using Qwen2.5-VL-7B-Instruct as the baseline model, we first fine-tune it via SFT on the step-by-step reasoning data of our proposed ChartRQA-SFT. This initial stage equips the model with the fundamental capability to decompose complex tasks into fine-grained subtasks. Our ablation studies demonstrate that this preliminary SFT stage on CoT data is critical, as it yields significantly better performance than applying RL from scratch. The model is trained using a standard autoregressive language modeling objective with negative log-likelihood loss.
@@ 3.3 Chart-RFT
After the Chart-COT stage, while the fine-tuned model demonstrates an enhanced ability to decompose complex questions, its performance on out-of-domain (OOD) tasks notably degrades. We hypothesize this is due to a distributional mismatch between ChartRQA-SFT with some simple chart understanding tasks, which harms its generalization ability. To address this degradation, we subsequently apply reinforcement fine-tuning (RFT) to generalize its reasoning capacity.
Group Relative Policy Optimization We adapt the Group Relative Policy Optimization (GRPO) algorithm Shao et al. (2024); Guo et al. (2025), which significantly conserves training resources by replacing the critic model with a baseline estimated from group scores. For each input $(x,y)$, the policy $\pi_{\theta}$ samples a group of $G$ candidate responses $\{o_{i}\}_{i=1}^{G}$.
$\mathcal{J}_{GRPO}(\theta)=$ $\mathbb{E}_{\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid x)\end{subarray}}$
$\biggl[\frac{1}{G}\sum_{i=1}^{G}\min\biggl(\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)}A_{i},$ (1)
$\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)},1-\varepsilon,1+\varepsilon\right)A_{i}\biggr)\biggr]$
where $\varepsilon$ is the hyperparameter, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the optimized model and the policy model respectively. The group-normalized advantage for the $i$-th response is:
@@
Reward Design For effective RFT, we follow the DeepSeek-R1 Shao et al. (2024) and adopt a rule-based reward function that consists of accuracy and format rewards to assess answer correctness and structural integrity, respectively.
  • Accuracy Reward. We employ distinct, type-specific reward functions to measure the correctness of model outputs. For numerical answers, we adopt the soft matching technique from Point-RFT Ni et al. (2025) with a relative error tolerance of $\pm$5%. For string-based answers, we utilize the edit distance as the reward signal.
  • Format Reward. The format reward is determined by a grammar-level regex parser that validates the structural integrity of outputs. It confirms two conditions: (1) the reasoning process is properly enclosed in <<think>> tags, and (2) the final answer is extractable from the designated <<answer>> tags.
@@
Model Name ChartQA CharXiv-RQ ChartQAPro ChartMuseum ChartRQA (single / multi)
Proprietary
GPT-4o 85.7 47.1 37.67 42.2 44.37 / 46.55
Gemini-1.5-Flash 79.0 33.9 42.96 31.1 -
Gemini-1.5-Pro 87.2 43.3 - 41.3 -
Gemini-2.5-Flash - - - - 59.12 / 59.17
Claude-3.5-Sonnet 90.8 60.2 43.58 54.4 52.79 / 56.05
GPT-4.1 86.8 56.7 - 48.4 57.88 / 59.30
Claude-3.7-Sonnet 86.1 64.2 - 60.3 55.04 / 57.87
General-domain Open-source
Phi-3.5-Vision 81.8 32.7 24.73 - 31.08 / 24.32
DeepSeek-VL2 86.0 - 16.28 - 23.15 / 20.29
InternVL3-8B 86.6 37.6 - 28.2 37.51 / 31.73
InternVL3-38B 89.2 46.4 - 32.1 46.09 / 38.36
Qwen2.5-VL-7B 87.3 42.5 36.61 26.8 44.59 / 40.57
R1-VL 83.9 32.0 35.11 21.0 27.33 / 21.72
VL-Rethinker 83.5 42.8 48.26 31.5 42.87 / 42.52
MMR1 83.7 40.5 45.72 30.9 47.27 / 46.55
Chart-domain
ChartLlama 69.66 14.2 - - -
TinyChart 83.60 8.3 13.25 12.5 6.75 / 6.11
ChartGemma 80.16 12.5 6.84 12.2 7.18 / 9.23
ChartReasoner 86.93 - 39.97 - -
BigCharts-R1 89.84 41.3 - - -
Bespoke-MiniChart 89.50 45.4 45.36 34.0 42.77 / 42.13
Chart-R1 (Ours) 91.04 46.2 44.48 34.8 52.09 / 49.93
Table 3: The main results on existing chart understanding and reasoning benchmarks. Our proposed Chart-R1 achieves superior performance among small-scale VLMs (<<20B) on the evaluation benchmarks. Bold denotes the best performances of open-source VLMs.